# Quantum Computing for Automotive Applications: From Algorithms to Applications

### v1

@@ 1 Introduction
This chapter is structured as follows: We begin by discussing the potential of quantum computing in the automotive industry in Section 2. Our analysis then focuses on four problem domains: optimization, simulation, materials science, and machine learning (sections 3-6). For each domain, we focus on three areas: (i) Near-term algorithms and techniques, such as variational quantum algorithms (VQAs) [13], aiming to exploit today’s quantum computers despite their limitations. (ii) Fault-tolerant quantum computing (FTQC) algorithms are expected to provide provable quantum advantages as hardware improves. (iii) Quantum-inspired approaches which deliver measurable benefits today. Section 7 discusses the importance of benchmarks for evaluating quantum and classical solutions. We summarize our results and discuss future work in section 8.
@@ 2 Automotive Applications Areas
Processes: Assembling vehicles is associated with many challenges related to the complexity of the product and the vast number of variants produced daily. Many optimization challenges arise in the context of decision-making and planning, involving strategic, tactical, and operational decisions. Examples of such problems are robotic path optimization [15], vehicle configurations [16], route optimization, placement & distribution problems [17, 18], line balancing, shift scheduling [19], and vehicle sequencing [3]. Many problems fall into the category of combinatorial optimization and are often NP-hard when solved brute-force, making them challenging for traditional computing methods. Quantum computing offers the potential to handle the computational demands of these problems.
@@
Numerical simulations are used to understand physical processes related to the vehicle, e. g., the airflow around a vehicle, which is government by the Navier-Stokes equation, or the behavior of electrical systems described by Maxwell’s equations. In practice, it is required to solve these equations over complex geometries under complex physical constraints. Quantum computing promises to solving these partial differential equations efficiently [22].
@@ 3 Optimization
In fault-tolerant quantum optimization algorithms, most approaches leverage Grover’s quantum search algorithm [26, 46]. An essential ingredient of Grover search is the so-called oracle, a function that identifies candidate solutions with desirable properties, e. g., feasibility. Krol et al. [19] construct an oracle for the shift scheduling problem, leading to a quadratic quantum speedup over a brute force search of the solution space. However, it remains unclear whether a quadratic speedup will be sufficient to achieve a quantum advantage in practice [47, 48].
@@
Several quantum-inspired optimization strategies have emerged as byproducts of the research in quantum optimization. Notably, the marriage of generative modeling and quantum-inspired tensor network methods resulted in the generator-enhanced optimization (GEO) framework [49]. In GEO, a quantum-inspired generative model is used to propose new candidate solutions by identifying patterns in high-quality solutions. The performance of GEO has since been analyzed on the production planning use case, where it has been shown to perform on par with other conventionally used solvers [50].
@@ 5 Materials Science and Quantum Chemistry
For material science and quantum chemistry, the starting point for quantum simulation is the Hamiltonian of the system. Usually denoted by $H$, the Hamiltonian describes the total energy of a given system, encodes the system’s physical properties, and defines its evolution via, for example, the Schrödinger equation for a closed quantum system. In general, the physical systems are described via spin or molecular models. Spin models describe the system of the particles via spin-spin interactions. For example, the Heisenberg spin model [104] can describe the electromagnetic properties of materials [105], and can readily be encoded into a quantum computer. Quantum systems can also be described by molecular models, which directly represent the electronic and nuclear interactions via electromagnetic forces among the atoms forming a molecule. This representation is commonly used in quantum chemistry. Ultimately, the molecular model can be mapped to a spin model via the second quantization and the Jordan-Wigner transformations [106].
@@ 8 Conclusion and Future Directions
General: This chapter examined the current state of quantum algorithms and their applications in the automotive industry. Table 1 summarizes the problem domain, applications, algorithms and case studies conducted using these algorithms. We categorize the algorithms into near-term, fault-tolerant, and quantum-inspired methods and map them to case studies across the automotive value chain.
@@
Quantum Hardware and Simulators: Table 2 provides a detailed overview of selected automotive case studies. The utilized hardware platforms range from D-Wave’s quantum annealers to Quantinuum and IonQ’s quantum processors and classical simulators. Additionally, the table captures the scale and complexity of the problems addressed, offering insight into quantum algorithms’ current capabilities and limitations as they are applied to real-world scenarios. In the following, we will discuss the key takeaways per problem domain, followed by the overarching trends.
@@
Notably, the data acquisition and problem modeling techniques developed in quantum projects can often be effectively transferred to classical solutions, ensuring that even with the current computational limitations of quantum hardware, significant results can be achieved.

### v2

@@ 1 Introduction
This chapter is structured as follows: We begin by discussing the potential of quantum computing in the automotive industry in Section 2. Our analysis then focuses on four problem domains: optimization, simulation, materials science, and machine learning (sections 3-6). For each domain, we focus on three areas: (i) Near-term algorithms and techniques, such as variational quantum algorithms (VQAs) [13], aiming to exploit today’s quantum computers despite their limitations. (ii) Fault-tolerant quantum computing (FTQC) algorithms are expected to provide long-term quantum advantages based on proven theoretical advantages; however, they also demand more reliable and scalable quantum hardware. (iii) Quantum-inspired approaches which deliver measurable benefits today. Section 7 discusses the importance of benchmarks for evaluating quantum and classical solutions. We summarize our results and discuss future work in section 8.
@@ 2 Automotive Applications Areas
Processes: Assembling vehicles is associated with many challenges related to the complexity of the product and the vast number of variants produced daily. Many optimization challenges arise in decision-making and planning, involving strategic, tactical, and operational decisions. Examples of such problems are robotic path optimization [15], vehicle configurations [16], route optimization, placement & distribution problems [17, 18], line balancing, shift scheduling [19], and vehicle sequencing [3]. Many problems fall into the category of combinatorial optimization and are often NP-hard when solved brute-force, making them challenging for traditional computing methods. Quantum computing offers the potential to handle the computational demands of these problems.
@@
Numerical simulations are used to understand physical processes related to the vehicle, e. g., the airflow around a vehicle, which is government by the Navier-Stokes equation, or the behavior of electrical systems described by Maxwell’s equations. In practice, it is required to solve these equations over complex geometries under complex physical constraints. Quantum computing promises to solve these partial differential equations efficiently [22].
@@ 3 Optimization
Most approaches in fault-tolerant quantum optimization algorithms leverage Grover’s quantum search algorithm [26, 46]. An essential ingredient of Grover search is the so-called oracle, a function that identifies candidate solutions with desirable properties, e. g., feasibility. Krol et al. [19] construct an oracle for the shift scheduling problem, leading to a quadratic quantum speedup over a brute force search of the solution space. However, it remains unclear whether a quadratic speedup will be sufficient to achieve a quantum advantage in practice [47, 48].
@@
Several quantum-inspired optimization strategies have emerged as byproducts of the research in quantum optimization. The marriage of generative modeling and quantum-inspired tensor network methods resulted in the generator-enhanced optimization (GEO) framework [49]. In GEO, a quantum-inspired generative model is used to propose new candidate solutions by identifying patterns in high-quality solutions. The performance of GEO has since been analyzed on the production planning use case, where it has been shown to perform on par with other conventionally used solvers [50].
@@ 5 Materials Science and Quantum Chemistry
For material science and quantum chemistry, the starting point for quantum simulation is the Hamiltonian of the system. Usually denoted by $H$, the Hamiltonian describes the total energy of a given system, encodes the system’s physical properties, and defines its evolution via, for example, the Schrödinger equation for a closed quantum system. In general, the physical systems are described via spin or molecular models. Spin models describe the system of the particles via spin-spin interactions. For example, the Heisenberg spin model [104] can describe the electromagnetic properties of materials [105] and can readily be encoded into a quantum computer. Quantum systems can also be described by molecular models, which directly represent the electronic and nuclear interactions via electromagnetic forces among the atoms forming a molecule. This representation is commonly used in quantum chemistry. Ultimately, the molecular model can be mapped to a spin model via the second quantization and the Jordan-Wigner transformations [106].
@@ 8 Conclusion and Future Directions
General: This chapter examined the current state of quantum algorithms and their applications in the automotive industry. Table 1 summarizes the problem domains, applications, algorithms and case studies conducted using these algorithms. We categorize the algorithms into near-term, fault-tolerant, and quantum-inspired methods. While much research has been focused near-term algorithms and application, increasingly fault-tolerant methods are investigated.
@@
Quantum Hardware and Simulators: The utilized hardware platforms range from D-Wave’s quantum annealers to Quantinuum and IonQ’s quantum processors and classical simulators. Additionally, the table captures the scale and complexity of the problems addressed, offering insight into quantum algorithms’ current capabilities and limitations as they are applied to real-world scenarios. In the following, we will discuss the key takeaways per problem domain, followed by the overarching trends.
@@
Notably, the data acquisition and problem modeling techniques developed in quantum projects can often be effectively transferred to classical solutions, ensuring that significant results can be achieved even with quantum hardware’s current computational limitations.